A K_{A}-BAND CHIRPED-PULSE FOURIER TRANSFORM MICROWAVE SPECTROMETER.

\underline{DANIEL P. ZALESKI}, JUSTIN L. NEILL, MATTHEW T. MUCKLE, AND BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904; P. BRANDON CARROLL AND SUSANNA L. WIDICUS WEAVER, Department of Chemistry, Emory University, 1515 Dickey Drive, Atlanta, GA 30322.

The design and performance of a new chirped-pulse Fourier transform microwave (CP-FTMW) spectrometer operating from 25-40 GHz will be discussed. A 10.5-3 GHz linear frequency sweep, generated by a 24 GS/s arbitrary waveform generator, is upconverted by a 23.00 GHz phase-locked oscillator, then fed into an active doubler to create a 25-40 GHz chirped pulse. After amplification with a 60-80 W pulsed traveling wave tube amplifier, the pulse is broadcast across a molecular beam chamber where it interacts with a molecular sample. The molecular FID signal is downconverted with the 23 GHz oscillator so that it can be digitized on a 50 GS/s oscilloscope with 16 GHz hardware bandwidth. The sensitivity and phase stability of this spectrometer is comparable to that of the previously reported 6.5-18.5 CP-FTMW spectrometer.^{a} On propyne (μ = 0.78 D), a single-shot signal to noise ratio of approximately 200:1 is observed on the J = 2 − 1 rotational transition at 34183 MHz when the full bandwidth is swept; optimal excitation is observed for this transition with a 250 MHz bandwidth sweep. The emission has a T_{2} lifetime of 4 μs. Early results from this spectrometer, particularly in the study of species of astrochemical interest, will be presented.

aG.G. Brown et al., Rev. Sci. Instrum. 79 (2008) 053103.